# Hallucination of Multimodal Large Language Models: A Survey

### v1

@@ 5.4.2. Post-hoc Correction
LogicCheckGPT (Wu et al., 2024) is a more recent self-revising-based hallucination mitigation method. Unlike Volcano (Lee et al., 2023), which revises the generated response with the help of general feedback, LogicCheckGPT delves into the logical consistency of MLLMs’ responses. Specifically, the approach can be formulated into two stages: the first stage involves inquiring attributes of objects, followed by inquiring objects based on attributes. Whether their responses can form a logical closed loop serves an indicator of object hallucination. If the ratio of closed loops to the total number of questions exceeds a certain threshold, rectify the hallucinated objects by prompting the MLLM.
@@ 6.2. Cross-modal Alignment and Consistency
The key challenge of multimodal hallucination is the cross-modal consistency issue. Ensuring that generated content remains consistent and contextually relevant to the input modality requires sophisticated techniques for capturing and modeling cross-modal relationships. The direction of cross-modal alignment encompasses both MLLMs training and hallucination evaluation. Regarding training, future research should explore methods for aligning representations between different modalities. Achieving this goal may involve designing more advanced architectures, introducing additional learning objectives (Jiang et al., 2023), or incorporating diverse supervision signals (Chen et al., 2023c). Regarding evaluation, cross-modal consistency checking has been a long-standing topic, ranging from multimodal understanding (Radford et al., 2021; Li et al., 2023b) to text-to-image generation (Cho et al., 2023; Chen et al., 2024a). Drawing on proven experiences from these domains to improve the assessment of MLLM hallucination, or unifying them into an overall framework, may be promising research directions.
@@ Acknowledgements.
This project is supported by Mike Zheng Shou’s Start-Up Grant from NUS.

### v2

@@ 5.4.1. Generation Intervention
The work of ProjectAway (Jiang et al., 2024b) explores a novel strategy that directly edits the latent image representation. Through analysis and experiment, they find the explored strategy can erase both real and hallucinated objects with high rates of removal. Therefore, this tool is utilized to detect hallucinations and reduce hallucination by erasing hallucinations from the VLM’s internal representations. Similar to editing latent space, Nullu (Yang et al., 2024) directly edits the model weights, based on an unsafe subspace, which is called HalluSpace in the paper. By orthogonalizing the model weights, input features will be projected into the Null space of the HalluSpace to reduce hallucination.
CausalMM (Zhou et al., 2024b) develops a causal inference framework that applies structural causal modeling to MLLMs, treating modality priors as a confounder between attention mechanisms and output. By applying counterfactual reasoning on the framework, it makes the output more aligned with multimodal inputs.
Some works propose various prompting strategies to mitigate hallucination. The work of  (Kim et al., 2024b) proposes to instruct the MLLM first think about some counterfactual contents. Then, at the actual generation step, the model should answer the user query while simultaneously avoid the counterfactual terms. VIC (Zheng et al., 2024b) designs a thinking before looking strategy to leverage the reasoning capability of the LLM, enabling MLLMs to anticipate and recognize patterns more efficiently by dynamically adjusting reasoning steps.
5.4.2. Visual Prompting
As many works (Huang et al., 2023b; Favero et al., 2024) have identified that one prominent cause of hallucination is weak or fading visual attention, one intuitive way to alleviate this issue is to explicitly highlight the objects-of-interests. An early exploration, SoM-LLaVA (Yan et al., 2024b), applies the Set-of-Mask prompting strategy (Yang et al., 2023) on MLLMs, demonstrating effectiveness on reducing hallucination.
The work of VAP (Zhang et al., 2025a) designs an solution: strategically crafted perturbation to visual inputs can redirect LVMs’ decision-making processes away from parametric biases without altering the original model’s architecture or mechanisms. It formulates an adversarial strategy to align the MLLM responses with visual content while reducing the impact of parametric knowledge bias. Eventually, the perturbation noise is optimized to alleviate parametric bias thus reducing hallucination. Essentially, this strategy can be regarded as a type of visual prompting.
Another work (Tomita et al., 2025) investigates the role of background information in reducing object hallucination from a visual prompting perspective.
5.4.3. RAG
Recently, in the realm of LLM, Retrieval-Augmented Generation (RAG) (Wang et al., 2020; Lewis et al., 2020) from external knowledge resources has shown promise in reducing language hallucinations. This retrieval augmentation serves as a flexible way to extend beyond the model’s inherent knowledge without the burden of extensive training costs. This technique has also been used to mitigate hallucination in MLLM. ARA (Qu et al., 2024a) adopts RAG by 1) decomposing the target object causing hallucination from the input image; 2) discerning the most effective retrieval technique and securing trustworthy results, considering the nature of multimodal; 3) initiating retrieval only during periods of low LVLM certainty and knowledge deficiency to reduce cost. A similar idea is further explored in a more recent work FilterRAG (RAG).
5.4.4. Ensembling
As a black box model, MLLM can be affected by even smaller perturbation in at the input model parameters, which may lead to hallucination. Therefore, it is intuitive to consider ensembling results from multiple inputs or models to obtain a more robust result.
RITUAL (Woo et al., 2024a) proposes to leverage random image transformations to complement the original image and enhance models’ robustness. The key idea is that by exposing the model to diverse visual transformations—such as changes in orientation, scale, and color—during decoding, it can better discern the true contents of the original image and reduce the likelihood of generating hallucinatory outputs.
MAD (Lin et al., 2024) addresses the hallucination problem from a model ensembling-style solution. Specifically, it adopts a multi-agent debate strategy by instantiating several MLLM as agents to conduct a multi-round debate. This cross-agent debate encourages divergent-thinking and promotes cognitive synergy.
Another ensembling-style solution is explored in MVP (Qu et al., 2024b), where a Multi-view-Multi-path strategy is proposed. It involves a multi-view information-seeking strategy that exhausts the perception of the image from varying dimensions: “top-down”, “bottom-up”, and “regular”. In addition, during the answer decoding stage, it further introduces multi-path reasoning for each information view by explicitly quantifying the certainty score of the potential answers and then aggregating the overall certainty among multiple paths.
The work of Dentist (Chang et al., 2024) argues that there are different types of queries that can be addressed by different strategies. Therefore, it proposes a unified framework that first classifies the queries, then performs different processes of hallucination mitigation based on the classification result.
@@ 5.4.5. Post-hoc Correction
LogicCheckGPT (Wu et al., 2024c) is a self-revising-based hallucination mitigation method. Unlike Volcano (Lee et al., 2023), which revises the generated response with the help of general feedback, LogicCheckGPT delves into the logical consistency of MLLMs’ responses. Specifically, the approach can be formulated into two stages: the first stage involves inquiring attributes of objects, followed by inquiring objects based on attributes. Whether their responses can form a logical closed loop serves an indicator of object hallucination. If the ratio of closed loops to the total number of questions exceeds a certain threshold, rectify the hallucinated objects by prompting the MLLM.
VFC (Ge et al., 2024) adopts an agent-style workflow to generate accurate, detailed caption. This is achieved by 1) proposal, where image-to-text captioning models propose multiple initial captions; 2) verification, where a large language model (LLM) utilizes tools such as object detection and VQA models to fact-check proposed captions; 3) captioning, where an LLM generates the final caption by summarizing caption proposals and the fact check verification results. The verification step can effectively mitigate hallucination. Dhall (Wu et al., 2024b) designs a holistic reasoning framework to mitigation hallucination, encompassing six reasoning modules.
@@ 6.2. Cross-modal Alignment and Consistency
The key challenge of multimodal hallucination is the cross-modal consistency issue. Ensuring that generated content remains consistent and contextually relevant to the input modality requires sophisticated techniques for capturing and modeling cross-modal relationships. The direction of cross-modal alignment encompasses both MLLMs training and hallucination evaluation. Regarding training, future research should explore methods for aligning representations between different modalities. Achieving this goal may involve designing more advanced architectures, introducing additional learning objectives (Jiang et al., 2023), or incorporating diverse supervision signals (Chen et al., 2023b). Regarding evaluation, cross-modal consistency checking has been a long-standing topic, ranging from multimodal understanding (Radford et al., 2021; Li et al., 2023b) to text-to-image generation (Cho et al., 2023; Chen et al., 2024c). Drawing on proven experiences from these domains to improve the assessment of MLLM hallucination, or unifying them into an overall framework, may be promising research directions.
@@ Acknowledgements.
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-030).